OpenStudy (anonymous):

help! State the domain and range for the function. f(x)=2csc(x/2))

myininaya (myininaya):

For the domain, you want to consider numbers such that sin(x/2) is not zero since we cannot divide by zero. So find when sin(x/2)=0 and those numbers will be the numbers we exclude from our domain.

myininaya (myininaya):

Now the range of y=csc(x) is (-inf,-1]U[1,inf) so you need to find the range of y=2*csc(x)...That just means our outputs for y=csc(x) will by multiplied by 2.

myininaya (myininaya):

First when is sin(x/2)=0?